OpenStudy (anonymous):

[8.06] Determine which polynomial is a perfect square trinomial. (3 points) 4x2 - 12x + 9 16x2 + 24x - 9 4a2 - 10a + 25 36b2 - 24b - 16 help?

OpenStudy (e.cociuba):

Let's start with the definition of a perfect square trinomial. Perfect square trinomial: An expression with three terms that is the square of a binomial. So now tell me. What does the square of a binomial look like?

OpenStudy (anonymous):

(x+a)^2 foil it out and it gives you the form for a perfect square :D

OpenStudy (anonymous):

(x+a)*(x+a)= x^2+ax+ax+a^2= x^2+2ax+a^2

OpenStudy (anonymous):

I know how to do this its just im unable to figure this problem out

OpenStudy (e.cociuba):

\[(a + b)^2 = a^2 + 2ab + b^2 or (a - b)^2 = a^2 - 2ab + b^2 \] These are the forms you must work with.

OpenStudy (e.cociuba):

Ok so lets start with: 4x^2 - 12x + 9 You are not "solving" anything, you are simplifying an expression. In fact you are trying to factor it.

OpenStudy (anonymous):

so in this one it would be 2x for first term and positive 3for last term

OpenStudy (anonymous):

i am not so sure what if we add another coefficien (bx+a)*(bx+a)= b^2x^2+bax+bax+a^2 (b^2)x^2+(2ba)x+a^2 maybe A B C Ax^2+ bx+ C

OpenStudy (anonymous):

so um what sign would go between 4x and 3? a negative or positive sign?

OpenStudy (e.cociuba):